# Faces of homogeneous cones and applications to homogeneous chordality

### v2

@@ 1 Introduction
where $q^{*}$ denotes the transpose of $q$, see [Pat00, Example 3.2.2], [BC75, Section 6]. That is, each face of $\mathcal{S}_{+}^{n}$ is linearly isomorphic to a smaller PSD cone and this isomorphism can be realized as an automorphism of $\mathcal{S}_{+}^{n}$ since the map $Q$ that takes $x$ to $qxq^{*}$ satisfies $Q(\mathcal{S}_{+}^{n})=\mathcal{S}_{+}^{n}$. The matrix $q$ does not come from thin air: the kernel of matrices in the relative interior of ${\mathcal{F}}$ is unique and correspond to some fixed subspace $V\subseteq\mathbb{R}^{n}$. With that, =${\mathcal{F}}=\{x\in\mathcal{S}_{+}^{n}\mid\ker x\supseteq V\}$ holds. Therefore, computing $q$ is an entirely constructive endeavour, as it is enough to pick any $x\in\mathrm{ri}\,{\mathcal{F}}$ and let $q$ be such that its rows are orthogonal and the last $n-r$ rows form an orthonormal basis for $\ker x$.
@@ 3. $(iii)$
    We discuss in detail the facial structure of PSD matrices with sparsity pattern given by a homogeneous chordal graph $G$, see Section 4. In particular we show that every face of $\mathcal{S}_{+}({G})$ is isomorphic to some face that arises by considering induced subgraphs of $G$ and this isomorphism can be realized by an automorphism of $\mathcal{S}_{+}({G})$, see Theorem 4.2. We will also prove analogous results to the dual cone of $\mathcal{S}_{+}({G})$ which corresponds to the PSD completable matrices determined by $G$. Then, we will check that when Algorithm 1 is specialized to the dual of $\mathcal{S}_{+}({G})$, it will allow us to compute PSD completions having certain desirable properties, see Theorem 4.3.
@@ 2 Preliminaries
[…]
Let ${\mathcal{K}}\subseteq\mathcal{E}$ be a convex cone. We say that ${\mathcal{K}}$ is pointed if $\mathrm{cl}\,{\mathcal{K}}\cap-\mathrm{cl}\,{\mathcal{K}}=\{0\}$ holds. Two convex cones ${\mathcal{K}}_{1},{\mathcal{K}}_{2}$ are said to be linearly isomorphic if there exists a linear bijection $A$ such that $A({\mathcal{K}}_{1})={\mathcal{K}}_{2}$ holds.
A convex cone ${\mathcal{F}}\subseteq{\mathcal{K}}$ is said to be a face of ${\mathcal{K}}$ if $x,y\in{\mathcal{K}}$ and $x+y\in{\mathcal{F}}$ implies that $x,y\in{\mathcal{F}}$. In this case, we write ${\mathcal{F}}\mathrel{\unlhd}{\mathcal{K}}$ and, by convention, we will only consider non-empty faces. A face ${\mathcal{F}}$ is said to be proper if ${\mathcal{F}}\neq{\mathcal{K}}$ and it is maximally proper if there is no proper face ^$\hat{\mathcal{F}}\mathrel{\unlhd}{\mathcal{K}}$ such that $^{\mathcal{F}}\subseteq\hat{\mathcal{F}}$ and $^{\mathcal{F}}\neq\hat{\mathcal{F}}$. An extreme ray of ${\mathcal{K}}$ is a face of the form =${\mathcal{F}}=\{\alpha x\mid\alpha\geq 0\}$ for some nonzero $x\in{\mathcal{K}}$. In this case, we say that $x$ generates an extreme ray of ${\mathcal{K}}$.
@@
In this proof we need the well-known fact that a face is exposed if and only if ${\mathcal{F}}^{\Delta\Delta}={\mathcal{F}}$ holds. In addition, all maximal proper faces are exposed, e.g., see [Tam85, Corollary 2.2]. With these facts in mind, suppose that ^$\hat{\mathcal{F}}\mathrel{\unlhd}{\mathcal{K}}^{*}$ is a maximally proper face containing ${\mathcal{F}}^{\Delta}$. As conjugacy inverts inclusion, we have ^$\hat{\mathcal{F}}^{\Delta}\mathrel{\unlhd}{\mathcal{F}}^{\Delta\Delta}={\mathcal{F}}$, so ^$\hat{\mathcal{F}}^{\Delta}=\{0\}$ or ^$\hat{\mathcal{F}}^{\Delta}={\mathcal{F}}$. In the former case we have ^=^$\hat{\mathcal{F}}=\hat{\mathcal{F}}^{\Delta\Delta}={\mathcal{K}}^{*}$, which contradicts the properness of $\hat{\mathcal{F}}$. So, we must be in the latter case and ^=^$\hat{\mathcal{F}}=\hat{\mathcal{F}}^{\Delta\Delta}={\mathcal{F}}^{\Delta}$ holds. ∎
@@
[…]
@@ 2.1 Homogeneous cones and T-algebras
Our goal in this paper is to describe the facial structure of homogeneous cones, which as far as we know, has not been described explicitly before. Nevertheless, there are previous works describing polynomials that are closely related to the boundary structure of homogeneous cones [Ish01, Nak14, Nak18, GIL24, Nak24]. A previous work by Ishi also contains important information on the action of the automorphism group on the boundary of the cone, see [Ish00]. We will revisit this point later in Section 3. In optimization, the study of homogeneous cones seems to have been initially motivated by self-concordant barriers and interior-point methods [G9̈6, GT98, Chu09], with other works focusing on geometric and representational aspects [TX01, Chu03, TT03, CT06].
In this work, we will use the theory of T-algebras of Vinberg [Vin63] as it seems to be the most natural for our purpose. Unfortunately, as of this writing, although the original russian text of [Vin63] is freely available on the internet, it is not completely trivial to get a copy of the English translation. Because of that, as we review T-algebras, we refer extensively to more accessible references in English such as [Chu03, CT06, Chu09, KTX12].
@@
$\mathcal{A}_{ij}\mathcal{A}_{kl}$ =$=\{0\}\qquad\text{if }j\neq k.$
@@
holds for $a,b\in\mathcal{A}$. A matrix algebra with involution is a matrix algebra equipped with a linear bijection $*:\mathcal{A}\to\mathcal{A}$ such that for
@@
With that, we have $(a^{*})_{ij}={a_{ji}^{*}}$, where “$a^{*}_{ji}$” should be read as $(a_{ji})^{*}$.
@@ 6. $({a}7)$
${\mathcal{K}(\mathcal{A})}$ is an open homogeneous convex cone in $\mathcal{H}(\mathcal{A})$. Conversely, let ${\mathcal{K}}$ be a pointed open homogeneous convex cone contained in a real finite dimensional space. Then, there is a T-algebra $\mathcal{A}$ for which ${\mathcal{K}(\mathcal{A})}$ is linearly isomorphic to ${\mathcal{K}}$.
It follows from Vinberg’s discussion in [Vin63] that the rank of a homogeneous cone is invariant under linear isomorphisms. We note this as a proposition with a short proof sketch using self-concordant barriers.
@@
Proof sketch.
Let ${\mathcal{K}}$ be a homogeneous cone of rank $r$. Güler and Tunçel showed that the smallest barrier parameter $\vartheta({\mathcal{K}})$ among all logarithmically homogenous self-concordant barriers of ${\mathcal{K}}$ satisfies $\vartheta({\mathcal{K}})=r$, see [GT98, Theorem 4.1]. Now, $\vartheta({\mathcal{K}})$ is invariant under linear isomorphisms which can be either verified directly or inferred from, say, [Nes18, Theorem 5.3.3]. ∎
@@
$\bar{\mathcal{A}}_{ij}\bar{\mathcal{A}}_{kl}$ =$=\{0\}\qquad\text{if }j\neq k.$
Defining $\bar{\rho}_{i}\coloneqq\rho_{n_{i}}$, $\bar{e}_{i}\coloneqq e_{n_{i}}$ and restricting the trace and involution functions to $\mathcal{A}_{I}$, it is straightforward (albeit tedious) to check that Axioms 1 through 7 still hold. For example, the restriction of the involution * is still an idempotent bijection from $\mathcal{A}_{I}$ to $\mathcal{A}_{I}$, because if $a_{ik}=a_{ki}=0$, then $a_{ik}^{*}=a_{ki}=0$ must also hold. Furthermore, we have
@@ 3.1 Unique Cholesky factorization in homogeneous cones
The following proposition was essentially proven by Gindikin [Gin92, Chapter 2, §1, Lemma 7], but we provide a different proof here from which we will be able to extract an algorithm.
@@
It remains to show the uniqueness of $t_{x}$. Suppose $x=tt^{*}$ for some proper $t\in\mathcal{T}_{+}$. As discussed in (3.3) and (3.4), the last column of $t$ is given by $u$ and $\hat{x}=(t-u)(t-u)^{*}=t_{\hat{x}}t_{\hat{x}}^{*}$ holds. In addition, since $t$ is proper, $t-u$ is a proper element in $\mathcal{T}_{+}(\mathcal{A}_{\{r\}})$, which implies $t-u=t_{\hat{x}}$ by the uniqueness of $t_{\hat{x}}$. Hence, we obtain $t=t_{\hat{x}}+u=t_{x}$, i.e., the uniqueness of $t_{x}$ follows. ∎
The proof of Proposition 3.2 shows that if $x\in\mathrm{cl}\,{\mathcal{K}(\mathcal{A})}$ and $x=tt^{*}$ holds for a proper $t\in\mathcal{T}_{+}$ then, the last column of $t$ is either zero (if $x_{rr}=0$) or is uniquely determined by (3.3). Denoting such a column by $u$ and the remaining $r-1$ columns of $t$ by $\hat{t}$, (3.4) implies that $\hat{t}$ satisfies
@@
Now, $\hat{t}$ is an element of $t\in\mathcal{T}_{+}(\mathcal{A}_{\{r\}})$, so again, the last column of $\hat{t}$ in $\mathcal{A}_{\{r\}}$ (i.e., the $(r-1)$-th column of $\hat{t}$ seen as an element of $\mathcal{A}$), is uniquely determined and is either zero (if $(x-uu^{*})_{r-1,r-1}=0$) or is given by (3.3) with $x-uu^{*}$ in place of $x$. This leads to an algorithm for computing a generalized Cholesky decomposition, see Algorithm 1. Note that Algorithm 1 can also be used to test membership in ${\mathcal{K}(\mathcal{A})}$: given some $x\in\mathcal{H}(\mathcal{A})$, we invoke Algorithm 1 with $x$ as input. Denoting by $t$ the output of Algorithm 1, we have $x\in{\mathcal{K}(\mathcal{A})}$ if and only if $x=tt^{*}$.
@@
0:  $x\in\mathrm{cl}\,{\mathcal{K}(\mathcal{A})}$
1:  $u\leftarrow 0$, $t\leftarrow 0$, $y\leftarrow x$
@@ 2. $(ii)$
    ${\mathcal{F}}$ is a homogeneous cone of rank $s<r$ and there is a triangular automorphism $Q_{u}$ of ${\mathcal{K}(\mathcal{A})}$ such that $Q_{u}({\mathcal{F}})=\mathrm{cl}\,{\mathcal{K}}(\mathcal{A}_{I})$ holds for some $I\subsetneq\{1,\ldots,r\}$. Furthermore, letting $x\in\mathrm{ri}\,{\mathcal{F}}$ and writing $x=tt^{*}$ for a proper $t\in\mathcal{T}_{+}(\mathcal{A})$, we have $I=\{i\mid\rho_{i}(t_{ii})=0\}$ and $s=r-|I|$.

### v3

@@ 1 Introduction
where $q^{*}$ denotes the transpose of $q$, see [Pat00, Example 3.2.2], [BC75, Section 6]. That is, each face of $\mathcal{S}_{+}^{n}$ is linearly isomorphic to a smaller PSD cone and this isomorphism can be realized as an automorphism of $\mathcal{S}_{+}^{n}$ since the map $Q$ that takes $x$ to $qxq^{*}$ satisfies $Q(\mathcal{S}_{+}^{n})=\mathcal{S}_{+}^{n}$. The matrix $q$ does not come from thin air: the kernel of matrices in the relative interior of $\mathcal{F}$} is unique and correspond to some fixed subspace $V\subseteq\mathbb{R}^{n}$. With that, ${\mathcal{F}}=\{x\in\mathcal{S}_{+}^{n}\mid\ker x\supseteq V\}$ holds. Therefore, computing $q$ is an entirely constructive endeavour, as it is enough to pick any $x\in\mathrm{ri}\,{\mathcal{F}}$ and let $q$ be such that its rows are orthogonal and the last nrn-r rows form an orthonormal basis for $\ker x$.
@@ 3. $(iii)$
    We discuss in detail the facial structure of PSD matrices with sparsity pattern given by a homogeneous chordal graph $G$, see Section 4. In particular we show that every face of $\mathcal{S}_{+}({G})$ is isomorphic to some face that arises by considering induced subgraphs of $G$ and this isomorphism can be realized by an automorphism of $\mathcal{S}_{+}({G})$, see Theorem 4.2. We will also prove analogous results to the dual cone of $\mathcal{S}_{+}({G})$ which corresponds to the PSD completable matrices determined by $G$. Then, we will check that when Algorithm 1 is specialized to the dual of $\mathcal{S}_{+}({G})$, it will allow us to compute PSD completions having certain desirable properties, see Theorem 4.4. In Remark 4.3 we will see that some of our results may fail if a graph is chordal but not homogeneous chordal.
@@ 2 Preliminaries
Let $\mathcal{E}$ be a finite dimensional Euclidean space equipped with an inner product $\langle\cdot,\cdot\rangle$ and corresponding induced norm $\lVert{\cdot}\rVert$. Let $C\subseteq\mathcal{E}$ be a convex set. We denote its relative interior and closure by $\mathrm{ri}\,C$ and $\mathrm{cl}\,C$, respectively. The smallest subspace of $\mathcal{E}$ containing $C$ is denoted by $\mathrm{span}\,C$.
Let $\mathcal{K}$$\subseteq\mathcal{E}$ be a convex cone. We say that $\mathcal{K}$} is pointed if $\mathrm{cl}\,{\mathcal{K}}\cap-\mathrm{cl}\,{\mathcal{K}}=\{0\}$ holds. $\mathcal{K}$} is said to be full-dimensional if $\dim{\mathcal{K}}=\dim\mathcal{E}$ or, equivalently, if $\mathcal{K}$} has non-empty interior. A pointed full-dimensional cone is said to be regular. Two convex cones ${\mathcal{K}}_{1},{\mathcal{K}}_{2}$ are said to be linearly isomorphic if there exists a linear bijection $A$ such that $A({\mathcal{K}}_{1})={\mathcal{K}}_{2}$ holds.
A convex cone ${\mathcal{F}}\subseteq{\mathcal{K}}$ is said to be a face of $\mathcal{K}$} if $x,y\in{\mathcal{K}}$ and $x+y\in{\mathcal{F}}$ implies that $x,y\in{\mathcal{F}}$. In this case, we write ${\mathcal{F}}\mathrel{\unlhd}{\mathcal{K}}$ and, by convention, we will only consider non-empty faces. A face $\mathcal{F}$} is said to be proper if ${\mathcal{F}}\neq{\mathcal{K}}$ and it is maximally proper if there is no proper face $\hat{\mathcal{F}}\mathrel{\unlhd}{\mathcal{K}}$ such that $^{\mathcal{F}}\subseteq\hat{\mathcal{F}}$ and $^{\mathcal{F}}\neq\hat{\mathcal{F}}$. An extreme ray of $\mathcal{K}$} is a face of the form ${\mathcal{F}}=\{\alpha x\mid\alpha\geq 0\}$ for some nonzero $x\in{\mathcal{K}}$. In this case, we say that $x$ generates an extreme ray of $\mathcal{K}$}.
@@
In this proof we need the well-known fact that a face is exposed if and only if ${\mathcal{F}}^{\Delta\Delta}={\mathcal{F}}$ holds. In addition, all maximal proper faces are exposed, e.g., see [Tam85, Corollary 2.2]. With these facts in mind, suppose that $\hat{\mathcal{F}}\mathrel{\unlhd}{\mathcal{K}}^{*}$ is a maximally proper face containing ${\mathcal{F}}^{\Delta}$. As conjugacy inverts inclusion, we have $\hat{\mathcal{F}}^{\Delta}\mathrel{\unlhd}{\mathcal{F}}^{\Delta\Delta}={\mathcal{F}}$, so $\hat{\mathcal{F}}^{\Delta}=\{0\}$ or $\hat{\mathcal{F}}^{\Delta}={\mathcal{F}}$. In the former case we have $\hat{\mathcal{F}}=\hat{\mathcal{F}}^{\Delta\Delta}={\mathcal{K}}^{*}$, which contradicts the properness of $\hat{\mathcal{F}}$. So, we must be in the latter case and $\hat{\mathcal{F}}=\hat{\mathcal{F}}^{\Delta\Delta}={\mathcal{F}}^{\Delta}$ holds. ∎
@@
If the projection $\mathcal{P}$ can be taken to be self-adjoint with respect to $\langle\cdot,\cdot\rangle$, then $\mathcal{F}$} is said to be orthogonally projectionally exposed (or o.p.-exposed). Clearly, this is a notion that depends on the choice of inner product, so to emphasize this choice, we will sometimes say that $\mathcal{F}$} is o.p.-exposed under $\langle\cdot,\cdot\rangle$. If all faces of $\mathcal{K}$} are o.p.-exposed with respect the same inner product, then $\mathcal{K}$} is said to be o.p.-exposed. Being o.p.-exposed is a more restrictive property and, for example, an o.p.-exposed polyhedral cone must be linearly isomorphic to a nonnegative orthant, see [ST90, Theorem 3.7 and pg. 233] and [BLP87]. That said, all symmetric cones are o.p.-exposed under an appropriate inner product [Lou21, Proposition 33].
@@ 2.1 Homogeneous cones and T-algebras
Our goal in this paper is to describe the facial structure of homogeneous cones, which as far as we know, has not been described explicitly before. Nevertheless, there are previous works describing polynomials that are closely related to the boundary structure of homogeneous cones [Ish01, Nak14, Nak18, GIL24, Nak24]. A previous work by Ishi also contains important information on the action of the automorphism group on the boundary of the cone, see [Ish00]. We will revisit this point later in Section 3. In optimization, the study of homogeneous cones seems to have been initially motivated by self-concordant barriers and interior-point methods [Gül96, GT98, Chu09], with other works focusing on geometric and representational aspects [TX01, Chu03, TT03, CT06].
In this work, we will use the theory of T-algebras of Vinberg [Vin63] as it seems to be the most natural for our purpose. We will later justify this choice and discuss related approaches in Section 3.4. Unfortunately, as of this writing, although the original russian text of [Vin63] is freely available on the internet, it is not completely trivial to get a copy of the English translation. Because of that, as we review T-algebras, we refer extensively to more accessible references in English such as [Chu03, CT06, Chu09, KTX12].
@@
$\mathcal{A}_{ij}\mathcal{A}_{kl}$ $=\{0\}\qquad\text{if }j\neq k.$
@@
holds for $a,b\in\mathcal{A}$. A matrix algebra with involution is a matrix algebra equipped with a linear bijection $\mathcal{A}\to\mathcal{A}$ such that
@@
With that, we have $(a^{*})_{ij}={a_{ji}^{*}}$, where $a^{*}_{ji}$” should be read as $(a_{ji})^{*}$.
@@
$\mathcal{K}(\mathcal{A})$} is an open homogeneous convex cone in $\mathcal{H}(\mathcal{A})$. Conversely, let $\mathcal{K}$} be a pointed open homogeneous convex cone contained in a real finite dimensional space. Then, there is a (up to isomorphism) unique T-algebra $\mathcal{A}$ for which $\mathcal{K}(\mathcal{A})$} is linearly isomorphic to $\mathcal{K}$}.
Let $\mathcal{K}$$\subseteq\mathcal{E}$ be a pointed homogeneous convex cone. By restricting $\mathcal{E}$ to $\mathrm{span}\,{\mathcal{K}}$, we may assume that $\mathcal{K}$} has non-empty interior. Then, we define the homogeneous cone rank of $\mathcal{K}$} as the rank of the corresponding T-algebra given in Theorem 2.3 for the interior of $\mathcal{K}$}. This is well-defined because isomorphic T-algebras must have the same rank, see [Vin63, Definition 5]. For a non-pointed homogeneous closed convex cone $\mathcal{K}$}, we may similarly define rank by looking at the rank of its pointed component ${\mathcal{K}}\cap L^{\perp}$, where $L\coloneqq{\mathcal{K}}\cap-{\mathcal{K}}$.
It follows from Vinberg’s discussion in [Vin63] that the rank of a homogeneous cone is invariant under linear isomorphisms. We note this as a proposition.
@@
We also note in passing that Güler and Tunçel showed that for a homogeneous pointed closed convex cone $\mathcal{K}$} with non-empty interior, the smallest barrier parameter $\vartheta({\mathcal{K}})$ among all logarithmically homogenous self-concordant barriers of $\mathcal{K}$} satisfies $\vartheta({\mathcal{K}})=r$, see [GT98, Theorem 4.1] and the discussion in [Chu09, Section 3.1]. Now, $\vartheta({\mathcal{K}})$ is also invariant under linear isomorphisms which can be either verified directly or inferred from, say, [Nes18, Theorem 5.3.3].
@@
$\bar{\mathcal{A}}_{ij}\bar{\mathcal{A}}_{kl}$ $=\{0\}\qquad\text{if }j\neq k.$
Defining $\bar{\rho}_{i}\coloneqq\rho_{n_{i}}$, $\bar{e}_{i}\coloneqq e_{n_{i}}$ and restricting the trace and involution functions to $\mathcal{A}_{I}$, it is straightforward (albeit tedious) to check that Axioms 1 through 7 still hold. For example, the restriction of the involution * is still an involutive bijection from $\mathcal{A}_{I}$ to $\mathcal{A}_{I}$, because if $a_{ik}=a_{ki}=0$, then $a_{ik}^{*}=a_{ki}^{*}=0$ must also hold. Furthermore, we have
@@ 3.1 Unique Cholesky factorization in homogeneous cones
The following proposition was essentially proven by Gindikin [Gin92, Chapter 2, §1, Lemma 7]. Chua also provides a similar result in [Chu06, Lemma 17] by first showing the analogous result for positive semidefinite matrices and making use of matrix realizations of homogeneous cones. Here, we provide a different proof from which we will be able to extract a factorization algorithm.
@@
It remains to show the uniqueness of $t_{x}$. Suppose $x=tt^{*}$ for some proper $t\in\mathcal{T}_{+}$. As discussed in (3.3) and (3.4), the last column of $t$ is given by $u$ and $\hat{x}=(t-u)(t-u)^{*}=t_{\hat{x}}t_{\hat{x}}^{*}$ holds. In addition, since $t$ is proper, tut-u is a proper element in $\mathcal{T}_{+}(\mathcal{A}_{\{r\}})$, which implies $t-u=t_{\hat{x}}$ by the uniqueness of $t_{\hat{x}}$. Hence, we obtain $t=t_{\hat{x}}+u=t_{x}$, i.e., the uniqueness of $t_{x}$ follows. ∎
The proof of Proposition 3.2 shows that if $x\in\mathrm{cl}\,{\mathcal{K}(\mathcal{A})}$ and $x=tt^{*}$ holds for a proper $t\in\mathcal{T}_{+}$ then, the last column of $t$ is either zero (if $x_{rr}=0$) or is uniquely determined by (3.3). Denoting such a column by $u$ and the remaining r1r-1 columns of $t$ by $\hat{t}$, (3.4) implies that $\hat{t}$ satisfies
@@
Now, $\hat{t}$ is an element of $t\in\mathcal{T}_{+}(\mathcal{A}_{\{r\}})$, so again, the last column of $\hat{t}$ in $\mathcal{A}_{\{r\}}$ (i.e., the (r1)(r-1)-th column of $\hat{t}$ seen as an element of $\mathcal{A}$), is uniquely determined and is either zero (if $(x-uu^{*})_{r-1,r-1}=0$) or is given by (3.3) with $x-uu^{*}$ in place of $x$. This leads to an algorithm for computing a generalized Cholesky decomposition, see Algorithm 1. Note that Algorithm 1 can also be used to test membership in $\mathcal{K}(\mathcal{A})$}: given some $x\in\mathcal{H}(\mathcal{A})$, we invoke Algorithm 1 with $x$ as input. Denoting by $t$ the output of Algorithm 1, we have $x\in{\mathcal{K}(\mathcal{A})}$ if and only if $x=tt^{*}$.
@@
[…]
@@ 2. $(ii)$
    $\mathcal{F}$} is a homogeneous cone of rank $s<r$ and there is a triangular automorphism $Q_{u}$ of $\mathcal{K}(\mathcal{A})$} such that $Q_{u}({\mathcal{F}})=\mathrm{cl}\,{\mathcal{K}}(\mathcal{A}_{I})$ holds for some nonempty $I\subseteq\{1,\ldots,r\}$. Furthermore, letting $x\in\mathrm{ri}\,{\mathcal{F}}$ and writing $x=tt^{*}$ for a proper $t\in\mathcal{T}_{+}(\mathcal{A})$, we have $I=\{i\mid\rho_{i}(t_{ii})=0\}$ and s=r|I|s=r-|I|.